An executive invests $24,000, some at 6% and some at 5% annual interest. If he receives an annual return of $1,380, how much is invested at each rate?
Answer by hamsanash1981@gmail.com(151)

Let the amount invested at 6% and 5% annual interest be x and y. the number of years be 1.
the x + Y = 24000 ---- 1
and simple interest for x is (x*1*6)/100 => 6X/100
simple interest for y is (y*1*5)/100 => 5y/100
Given that he recieves an annual returns of $1,380, then
6x/100 + 5y/100 = 1380 --- 2
multipling equation 2 by 100
6x+ 5y = 138000 --- 3
Solving equ 1 and 3
1*6 6x + 6y = 144000
6x + 5y = 138000
----------------
y = 6000
then x = 24000 - 6000 => x = 18000
therefore 18000 is invested at 6% and 6000 is invested at 5%.